# Bringing Diversity from Diffusion Models to Semantic-Guided Face Asset Generation

### v1

@@ 3.1.3. Texture Data Normalization
The goal of this step is to remove lighting and other external visual effects, such as makeup, facial hair, glasses, and shadows, from the UV textures created in the last step. The normalized texture can be considered as clean albedo map that is a part of the output 3D face asset. While FFHQ-UV (Bai et al., 2023) attempts to normalize this data into clean albedo in UV form, it struggles to fully preserve identities and attributes. We pose texture normalization in this context as a domain transfer problem from the source domain, which consists of synthetic textures under different lighting as shown in Figure 3 (b), to the target domain, which consists of clean albedo as shown in Figure 3 (a). Since the target domain contains only collected 200 identities while the source domain containing around 44K identities, a convolutional image-to-image translation model directly trained on the whole image would face a high risk of overfitting. The following text describes several designs to mitigate this issue.
First, we introduce an assumption that lighting affects appearance of the face approximately uniformly in a reasonably-sized patch. Therefore we divide an input image into a grid of patches and assign a spatially varying factor to each grid corners to account for factors that are independent of albedo. is bilinearly interpolated in the interior of each patch. Under this formulation, image translation is formulated as a function parameterized by an mlp network that takes the image as input and translates each pixel independently based on the spatially-varying factor . However, the image translation is preceded by a patch parameter estimation network, a convolutional network that takes the image as input and computes at patch corners. Effects such as subsurface scattering and inaccuracies due to the non-physical nature of the input images can also be handled by the spatial variation of and do not need to be considered separately.
@@
The loss function for training the texture normalization model is constructed as:
where be the set of scanned albedo texture. During training, we additionally augment the input with albedo images that do not contain lighting and optimize the network to produce identical output to the input in this situation.
The overall training procedure is shown in Figure 6. As commonly done in unpaired image-to-image translation, a patch-based discriminator is employed to ensure that the translated image is in the target domain, trained with the binary cross entropy loss. For adversarial training the distribution of skin color of the normalized synthesized textures is optimized to match that of the scanned textures, so when normalizing a synthesized texture, the skin color is drawn randomly from the set of skin colors of the scanned textures. Let be the discriminator and the set of synthetic textures. The Discriminator’s loss and the normalization model’s adversarial loss are:
@@ 3.2. Base Geometry and Albedo Generation
With the prepared dataset, we train a geometry and albedo texture generation network in which each attribute of interest can be adjusted independently. In addition, the network is trained to be capable of image inversion and can disentangle identity information from attribute controls, such that it can be used as an attribute editor for a fixed identity. Our method follows the two-step approach in (Xiang et al., 2021). The overall procedure is shown in Figure 7.
@@
In the first step, a GAN with an autoencoder , a generator and a discriminator is designed to disentangle unlabeled information from the labels (in our case, the gender, age and ethnicity attributes). is first used to extracts unlabeled information from the images and discards labeled information . The code computed by the encoder therefore describes information that is independent of the face attributes (e.g. identity information).
@@
Through experiments, we noticed that although concatenating the one-hot attribute labels to the fully connected part of the discriminator worked as intended, incorporating the unlabeled code in the same way did not give satisfactory results. We speculate the difficulty lies in the observation that the unlabeled code is much longer than the attribute labels and has a strong spatial structure which the discriminator had to learn from scratch. Therefore, we help the discriminator by generating an image from the label and code it receives using the first-step generator and concatenating its output with the input image, so that the spatial structure is provided by the condition image and need not be re-learned. The detailed architecture of our discriminator is shown in the inset in Figure 7. remains frozen and is not updated along with other parameters of the discriminator.
3.3. Asset Refinement
A complete set of textures for physically-based rendering includes more than just albedo maps: it also comprises specular and displacement maps. However, our generated albedo and geometry, which are relatively coarse with 1K vertices and 1K resolution albedo maps, lack this information. In this section, we describe a refinement method that enhances our output by adding the missing assets and improving the resolution. This method includes three components.
1) Super-resolution: we train an super-resolution network (Chen et al., 2023b) to increase the 1K albedo to a 4K resolution with high-frequency details. This network not only enhances the clarity of the albedo but also recovers skin details that are lost in 1K resolution.
2) Specular and displacement maps: We train a translation network (Liang et al., 2021) that infers plausible specular and displacement maps from the 4K albedo obtained from the super-resolution network. Our model is trained using our high-quality scanned dataset. Specifically, we convert the albedo map into Lab color space and use only the L channel as input to train the displacement map network. The L channel encodes structure and illumination, while the ab channels encode color. This approach eliminates the influence of skin color, as the high-frequency geometry interpreted in displacement correlates only with shape, not skin color.
3) Secondary assets: As the final step, we integrate the missing secondary assets (such as the eyeball, teeth, and gum) and predefined Blendshapes into the obtained base mesh to prepare it for animation.
@@ 4.1. Implementation Details
Portraits Generation.
To generate the portraits Figure 2, we use the pretrained Stable Diffusion v1.5 (Rombach et al., 2022) as base LDM model and use the ControlNet checkpoint pre-trained on normal maps (Zhang, 2023) to add conditional controls for our frontal view normal map. The sampling method is DDIM with 40 time steps. Classifier-Free Guidance (Ho and Salimans, 2022) scale is 9.0. The output portraits resolution from the LDM model is 1024 1024.
@@
In this paper, we perform single view 3D face reconstruction to update geometry. We keep the camera fixed during normal map rendering, which is used as input for the diffusion model. The resulting portraits thus have a fixed camera pose, . To train the model, we rendered around synthetic frontal view portraits under different illuminations by using assets in our 3D high quality database. We combined these with captured multi-view real data. A variant of ReFA (Liu et al., 2022) was used as the baseline to train the model, with the input modified from multi-view to single view. The camera optimization was kept fixed during training, as we had the ground truth , so in each network step only position map get updated. The texture inference step is also skipped. All the training of the reconstruction model is performed on NVIDIA A100 graphics cards. The network parameters are randomly initialized and are trained using the Adam optimizer for 120,000 iterations with a learning rate set to . For the recurrent face geometry optimizer, we set the inference step to , the grid resolution to , the search radius to . An example of the refinement to the initial geometry is showed in Figure 8, the refined geometry (c) more accurately matches the portrait (b) compared to the initial geometry (a), with deeper-set eyes that are correctly positioned, a mouth that is properly sized and shaped, more correctly aligned cheekbones, and a chin that reflects the correct depth and contour, resulting in an overall head shape that aligns more closely with (b).
@@
To ensure that the discriminator focuses on only the lighting, its capacity is purposefully limited: it consists of an MLP with 6 layers and 64 features in the hidden layers. It takes 16 16 patches as input and first downsamples them to 4 4 so that the number of input features is .
@@
Our base geometry and albedo generation network generates at 0.014 seconds per subject. For 1k to 4K albedo upsampling, we trained a network (Chen et al., 2023b) on our high-quality dataset, which takes 53 seconds to complete in average. We trained two separate translation networks (Liang et al., 2021) for the specular and displacement maps and the whole translation takes 71 seconds in average. Utilizing these three types of networks, production-quality face assets for an identity can be created in less than 3 minutes on one Nvidia A6000 GPU.
@@ 4.2. Results
Qualitative Results of Generation.
Figure 9 shows full assets of 16 identities along with rendered images, created from 8 distinct semantic labels. For each row, the geometry, diffuse albedo map, specular map, and displacement map are created from a set of attributes and rendered under three different lighting conditions. In each row, we display two randomly generated results that follow the same input semantic attributes, yet represent different identities. These results illustrate the diversity, quality, and effective semantic control of our model.
Ours Diffusion
@@
Figure 4 showcases the direct output of our normalization network. It transfers the UV-space texture with illumination baked-in into albedo textures while preserving the identity of the input texture. We can further edit the skin tone as shown in Figure 14. For a quantitative evaluation, we calculate the Brightness Symmetry Error (BS Error) introduced by (Bai et al., 2023) of different datasets as shown in Table 1. For the Facescape data, we processed it using a method similar to (Li et al., 2020b) to obtain texture data with the same topology. For FFHQ, we employed a 3DMM fitting method to obtain partial textures from the images and then applied the blending described in Section 3.1.2 to complete the textures around the central facial area.
| Method | Facescape* | Light Stage* | FFHQ | Ours |
| BS Error | 15.2595 | 4.8279 | 28.0723 | 5.8109 |
@@
We also evaluate the accuracy of attribute control quantitatively. For this purpose, we train a classifier for age, gender and ethnicity, independent from the generator training process. The classifier is trained on 85% of normalized data and tested on the other 15%. We then generate a large number of samples using random codes and attribute labels, and classify these generated samples using the classifier. If the generator controls the attributes accurately, the performance of the classifier should be similar on the test data and on the generated samples. The accuracy (average of all classes) is shown in Table 2, and the confusion matrix is shown in Figure 15. The performance of the classifier on the dataset and the generated samples are similar, and where they differ, the accuracy is generally higher on the generated samples.
@@
South Asian Female Age 28 Germanic Unisex Age 40
Figure 14 shows independent control of skin color. In particular, we show the variation of skin color within two subjects, conditioned on their respective ethnicity. We calculate the skin color from each unnormalized texture using the method in (Thong et al., 2023) and for each ethnicity the distribution of skin colors is modeled by a normal distribution. The figure presents five skin color samples of each subject, arranged from bright to dark.
@@
| race | gender | age | |
| dataset | |||
| generated | |||
| difference |
Inversion and Editing.
Figure 12 shows image inversion and editing. Images are first projected into the latent space of the generator by finding the unlabeled code and labeled attributes that generates the closest image using optimization. Edited images are then generated by modifying the labeled attributes while keeping the unlabeled code unchanged.
Additionally, since all generated facial geometries share the same topology with our high-quality dataset, generic facial geometric attributes editing such as add aging features Figure 11 can be done with some predefined geometry offsets, and blendshapes can be applied to animate the facial assets. For animation examples, please refer to the video.
Comparison.
We access the accuracy of attribute control of our generator using CLIP score. We first build a text description of the generated image data for our semantic attributes. We utilize the same ethnicity, gender, and age groups as the attributes we defined for generation to construct descriptors. Following the strategy of (Zhang et al., 2023a), we form text input with the phrase “the realistic face of [DESCRIPTOR]”. After acquiring 5 sets of PBR assets for each descriptor, we render 2500 test images. To calculate the score of Describe3D (Wu et al., 2023), we synthesize facial images by generating the descriptors with the rules mentioned in their work and used the same phrase as the input into CILP model. Subsequently, we compute the average CLIP score from the ViT-B/16 and ViT-L/14 models. The result, shown in Table 3, highlights effectiveness of our method in achieving semantic coherence between the generated images and their corresponding text descriptions.
| Method | CLIP Score |
| DreamFace (Zhang et al., 2023a) | 0.291 ± 0.020 |
| Describe3D (Wu et al., 2023) | 0.284 ± 0.054 |
| UltrAvatar (Zhou et al., 2024) | 0.301 ± 0.023 |
| Our Method | 0.316 ± 0.042 |
Figure 16 compares our system to DreamFace (Zhang et al., 2023a). Our system generates real albedo, which reflects true skin color, whereas DreamFace only produces texture under evenly distributed illumination. Also, our generated face model exhibits greater diversity and more closely aligns with the user description as the training dataset of DreamFace mainly consists of Asian people. Compared to general text-to-3D methods like DreamFusion (Poole et al., 2022), LucidDreamer (Liang et al., 2023) and TRELLIS (Xiang et al., 2024), our method produces avatars with impressive detail and production-level textures. DreamFusion creates clear structures, but misses finer skin details. LucidDreamer and TRELLIS outputs vivid, detailed and realistic faces, however with strong artifacts.
Figure 17 shows the results of texture normalization compared to FFHQ-UV (Bai et al., 2023). Our normalization results are closer to the target albedo domain, with pure skin color and no lighting baked-in (e.g., highlights). However, the normalized UV generated by FFHQ-UV still retains some of the original illumination, which cannot be used for high-end PBR-base shader as albedo.
User Study
We conduct a user study for the output of avatar generation methods among our proposed method, DreamFace (Zhang et al., 2023a) and Describe3D (Wu et al., 2023). 87 participants rate images synthesized using the corresponding methods in two aspects: description consistency and photorealism. The outcomes in Figure 18 showcase our proposed method consistently outperforming others in terms of description consistency and photorealism.
Interactive Avatar Creation System.

### v2

@@ 3.1.3. Texture Data Normalization
The goal of this step is to remove lighting and other external visual effects, such as makeup, facial hair, glasses, and shadows, from the UV textures created in the last step. The normalized texture can be considered a clean albedo map that is a part of the output 3D face asset. FFHQ-UV (Bai et al., 2023) is the latest work that can normalize face texture into an albedo image, but we observe that it still struggles to fully preserve identities and facial attributes when applied to our synthesized portraits (see Table 2). We therefore propose a more effective normalization method, by posing texture normalization in this context as a domain transfer problem from the source domain, which consists of synthetic textures under different lighting as shown in Figure 3 (b), to the target domain, which consists of clean albedo as shown in Figure 3 (a). Since the target domain contains only 200 identities while the source domain contains around 44K identities, a convolutional image-to-image translation model directly trained on the whole image would face a high risk of overfitting. The following text describes several designs to mitigate this issue.
First, we assume that lighting affects the appearance of the face approximately uniformly in a reasonably-sized patch. Therefore we divide an input image into a grid of patches and assign a spatially varying factor to each grid corner to account for factors that are independent of albedo. is bilinearly interpolated in the interior of each patch. Under this formulation, image translation is formulated as a function parameterized by an MLP network that takes the image as input and translates each pixel independently based on the spatially-varying factor . However, the image translation is preceded by a patch parameter estimation network, a convolutional network that takes the image as input and computes at patch corners. Effects such as subsurface scattering and inaccuracies due to the non-physical nature of the input images can also be handled by the spatial variation of and do not need to be considered separately.
@@
The loss function for training the texture normalization model is constructed as where is the set of scanned albedo textures. During training, we additionally augment the input with albedo images that do not contain lighting and optimize the network to produce identical output to the input in this situation.
The overall training procedure is shown in Figure 6. As commonly done in unpaired image-to-image translation, a patch-based discriminator is employed to ensure that the translated image is in the target domain, trained with the binary cross entropy loss. For adversarial training, the distribution of skin color of the normalized synthesized textures is optimized to match that of the scanned textures, so when normalizing a synthesized texture, the skin color is drawn randomly from the set of skin colors of the scanned textures. Let be the discriminator and the set of synthetic textures. The Discriminator’s loss and the normalization model’s adversarial loss are:
@@ 3.2. Base Geometry and Albedo Generation
As mentioned in the introduction, we aim to learn a geometry and albedo texture generation model that can 1) effectively disentangle labeled attributes and unlabeled information (e.g. identity), 2) provide an easy way to perform image inversion, and 3) generate efficiently. With the prepared dataset (geometry, albedo texture with labeled attributes), we propose a two-step generative model, extending from the two-step approach in (Xiang et al., 2021), that satisfies all previous requirements, and we show in Section 4.2.3 that the design is more effective compared to the common practices of generation directly conditioning on the given attributes. The overall procedure is shown in Figure 7.
@@
In the first step, a GAN with an autoencoder , a generator and a discriminator is designed to disentangle unlabeled information from the labels (in our case, the gender, age and ethnicity attributes). is first used to extract unlabeled information from the images and discards labeled information. The code computed by the encoder therefore describes information that is independent of the face attributes (e.g. identity information).
@@
Through experiments, we noticed that although concatenating the one-hot attribute labels to the fully connected part of the discriminator worked as intended, incorporating the unlabeled code in the same way did not give satisfactory results. We speculate that the difficulty lies in the observation that the unlabeled code is much longer than the attribute labels and has a strong spatial structure which the discriminator had to learn from scratch. Therefore, we help the discriminator by generating an image from the label and code it receives using the first-step generator and concatenating its output with the input image, so that the spatial structure is provided by the condition image and need not be re-learned. The detailed architecture of our discriminator is shown in the inset in Figure 7. remains frozen and is not updated along with other parameters of the discriminator.
@@ 4.1. Implementation Details
Synthetic Training Data Generation.
To generate the portraits Figure 2, we use the pre-trained Stable Diffusion v1.5 (Rombach et al., 2022) as the base LDM model and use the ControlNet checkpoint pre-trained on normal maps (Zhang, 2023) to add conditional controls for our frontal view normal map. The sampling method is DDIM with 40 time steps. Classifier-Free Guidance (Ho and Salimans, 2022) scale is 9.0. The output portrait resolution from the LDM model is 1024 1024.
@@
In this paper, we perform single-view 3D face reconstruction to predict the 3D geometry of the generated portraits. To train the model, we rendered around synthetic frontal view portraits under different illuminations by using assets in our 3D high-quality database. We combined these with captured multi-view real data. A variant of ReFA (Liu et al., 2022) was used as the baseline to train the model, where the network is modified to reconstruct from a single-view image. The camera optimization is kept fixed during training, as the ground truth pose is known. Therefore, in each network step, only the position map , which corresponds to the 3D geometry of the face, is updated.
The training of the reconstruction model is performed on NVIDIA A100 graphics cards. The network parameters are randomly initialized and are trained using the Adam optimizer for 120,000 iterations with a learning rate set to . For the recurrent face geometry optimizer, we set the inference step to , the grid resolution to , the search radius to .
An example of the refinement to the initial geometry is shown in Figure 8, the refined geometry (c) more accurately matches the portrait (b) compared to the initial geometry (a), with deeper-set eyes that are correctly positioned, a mouth that is properly sized and shaped, more correctly aligned cheekbones, and a chin that reflects the correct depth and contour, resulting in an overall head shape that aligns more closely with (b).
@@
To ensure that the discriminator focuses on only the lighting, its capacity is purposefully limited: it consists of an MLP with 6 layers and 64 features in the hidden layers. It takes patches as input and first downsamples them to so that the number of input features is .
@@
Asset Refinement
We apply an asset refinement process that augments resolution and supplements missing components required for physically-based rendering. This process consists of the following three stages: First, a super-resolution network (Chen et al., 2023b) is used to upsample the 1K albedo map to 4K, recovering high-frequency skin details. Second, a texture translation network (Liang et al., 2021) predicts corresponding specular and displacement maps from the 4K albedo. The network is trained on our high-quality scanned dataset. To generate displacement maps, the albedo is converted to Lab color space and only the L channel is used as input, isolating geometric features from skin color. We also applied a filter that smooths out the edge of the displacement maps. Finally, we integrate secondary assets such as eyeballs, teeth, and gums, along with predefined Blendshapes, into the base mesh to support animation and rendering.
@@
Our base geometry and albedo generation network generates results in 0.014 seconds per subject. For 1K to 4K albedo upsampling, we trained a network (Chen et al., 2023b) on our high-quality dataset, which takes 53 seconds to complete on average. We trained two separate translation networks (Liang et al., 2021) for the specular and displacement maps and the whole translation takes 71 seconds on average. Utilizing these three types of networks, production-quality face assets for an identity can be created in less than 3 minutes on one Nvidia A6000 GPU.
South Asian Female Age 28 Germanic Unisex Age 40
@@ 4.2. Results
4.2.1. Qualitative Results
Face Model Generation
Figure 10 shows full assets of 16 identities along with rendered images, created from 8 distinct semantic labels. For each row, two randomly generated results are shown (column 2-5, column 6-9) that follow the same input semantic attributes under different identities. For each subject, the geometry, diffuse albedo map, specular map, and displacement map are created from a set of attributes (1st column) and rendered under three different lighting conditions. These results illustrate the diversity, quality, and effective semantic control of our model.
@@ 4.2.1. Qualitative Results
The result of texture normalization is shown as part of Figure 4, where columns 3 and 4 show training data before and after normalization. Skin color is given in column 1. It transforms the UV-space texture with baked-in illumination into albedo textures while preserving the identity of the input texture. Figure 12 shows the effect of skin color control in texture normalization.
@@
Figure 14 shows independent control of skin color in the generator. When skin color is not specified, it is sampled randomly from the skin color distribution of the chosen ethnicity, which is a normal distribution fit to the skin color distribution in the unnormalized training data. The figure presents five skin color samples of each subject, from bright to dark in the range of their ethnicity, while keeping identity and facial features consistent.
Inversion and Editing.
Figures 15 and 16 shows image inversion and editing. Images are first projected into the latent space of the generator by finding the unlabeled code and labeled attributes that generates the closest image using optimization. Edited images are then generated by modifying the labeled attributes while keeping the unlabeled code unchanged.
Additionally, since all generated facial geometries share the same topology from our 3D asset dataset , generic facial geometric attributes editing such as add aging features Figure 13 can be done with some predefined geometry offsets, and blendshapes can be applied to animate the facial assets. For animation examples, please refer to the video.
4.2.2. Comparative Evaluation
Texture normalization.
For a quantitative evaluation, we can take a UV-space image rendered using scanned data, normalize it, and compare the result to the ground truth albedo by calculating the PSNR. We compare our method with FFHQ-UV (Bai et al., 2023). However, in FFHQ-UV, the facial geometry fitting and texture normalization are integrated, and it cannot be applied to UV textures directly. So, we render our scanned data and apply FFHQ-UV. The result is in Table 1.
In addition, on datasets without known ground truth we calculate the Brightness Symmetry Error (BS Error) introduced in FFHQ-UV, as shown in Table 2. The metric measures the dissimilarity of brightness between the left and right halves of the face, based on the assumption that without lighting the brightness of the two halves should be identical. For the Facescape data, we processed it using a method similar to (Li et al., 2020b) to obtain texture data with the same topology. For FFHQ, we employed a 3DMM fitting method to obtain partial textures from the images and then applied the blending described in Section 3.1.2 to complete the textures around the central facial area. For a qualitative evaluation, Figure 9 shows the results of texture normalization compared to FFHQ-UV (Bai et al., 2023). Our normalization results are closer to the target albedo domain, with pure skin color and no lighting baked-in (e.g., highlights). However, the normalized UV generated by FFHQ-UV still retains some of the original illumination, which cannot be used for high-end PBR-base shader as albedo.
| Method | FFHQ-UV | Ours |
| PSNR | 17.64 | 24.67 |
| Method | Facescape* | Light Stage* | FFHQ | FFHQ- UV** | Ours |
| BS Error | 15.2595 | 4.8279 | 28.0723 | 7.293 | 5.8109 |
Facial Attribute Control.
Figure 17 compares our system to DreamFace (Zhang et al., 2023a), a state-of-the-art face asset generation approach. Our system generates real albedo, which reflects true skin color, whereas DreamFace only produces texture under evenly distributed illumination. Also, our generated face model exhibits greater diversity and more closely aligns with the user description as the training dataset of DreamFace mainly consists of Asian people. Compared to general text-to-3D methods like DreamFusion (Poole et al., 2022), LucidDreamer (Liang et al., 2023) and TRELLIS (Xiang et al., 2024), our method produces avatars with impressive detail and production-level textures. DreamFusion creates clear structures, but misses finer skin details. LucidDreamer and TRELLIS output vivid, detailed and realistic faces, but with strong artifacts.
@@ 4.2.2. Comparative Evaluation
Ours Diffusion
For a quantitative comparison of attribute control of our generator to the other methods, we use CLIP score. We first build a text description of the generated image data for our semantic attributes. We utilize the same ethnicity, gender, and age groups as the attributes we defined for generation to construct descriptors. Following the strategy of (Zhang et al., 2023a), we form text input with the phrase “the realistic face of [DESCRIPTOR]”. After acquiring 5 sets of PBR assets for each descriptor, we render 2500 test images. To calculate the score of Describe3D (Wu et al., 2023), we synthesize facial images by generating the descriptors with the rules mentioned in their work and used the same phrase as input into the CLIP model. Subsequently, we compute the average CLIP score from the ViT-B/16 and ViT-L/14 models. The result, shown in Table 3, highlights the effectiveness of our method in achieving semantic coherence between the generated images and their corresponding text descriptions.
User Study
We conduct a user study to compare the outputs of avatar generation methods, including our proposed method, DreamFace (Zhang et al., 2023a) and Describe3D (Wu et al., 2023). 87 participants rate images synthesized using the corresponding methods in two aspects: description consistency and photorealism. The outcomes in Figure 20 show that our proposed method consistently outperforms the baseline methods in terms of description consistency and photorealism.
4.2.3. Ablation Study
Design of the Texture Normalization Network.
To verify our design for the normalization network, we compare it with the following variations: 1) “uniform patch parameter”: removing the spatial variability of patch parameter by pooling the output of the patch parameter network and broadcasting it to every location, 2) “no color control”: removing external color control and instead using the color calculated from the flat regions of the input image, 3) “Unet”: replacing the normalizer with a UNet, 4) “stride 1 conv net”: replacing the normalizer with a simple convolution network containing only stride-1 convolutions. The results are shown in Figure 21. The network with uniform patch parameters only applies a global color mapping. Due to the invariability of patch parameters it cannot handle strong local variations in lighting, as shown in row 2. Without color control, the quality is comparable, but the model loses its ability to incorporate color information if such information is available. In these examples, the race is known, and the color input is drawn randomly from the color distribution of the race, and the network without color control fails to produce race-accurate output skin color. The UNet enables long-range exchange of information and could modify facial features if that results in a better match between the output normalized texture and the high-quality scan. The scanned data contains a large portion of subjects with heavy eyebrows, and it can be observed that in row 4, the eyebrow appears darker and more expansive in the UNet output than the ground truth. In the stride-1 convnet, as each output pixel depends only on a small patch in the input, there is no coordination across the image, which can result in major failure cases as in row 1. Overall, the ablation study validated that our normalization design performs the best compared to the alternatives.
Effectiveness of Attribute Control.
We evaluate whether, for any given label, the distribution of the features in the images produced by our method closely match the real data distribution. For this purpose, we train a classifier for age, gender and ethnicity, independent from the generator training process. The classifier is trained on 85% of normalized data and tested on the other 15%. We then generate a large number of samples using random codes and attribute labels, and classify these generated samples using the classifier. If the generator controls the attributes accurately, the performance of the classifier should be similar on the test data and on the generated samples. The accuracy (average of all classes) is shown in Table 4, and the confusion matrix is shown in Figure 22. The performance of the classifier on the dataset and the generated samples are similar, and where they differ, the accuracy is generally higher on the generated samples.
| race | gender | age | |
| dataset | |||
| generated | |||
| difference |
@@ 4.2.3. Ablation Study
Identity Consistency in Attribute Editing.
A key demand for semantic editing of faces is the consistency in facial identity as the attributes of interest (age, gender and ethnicity in our case) are adjusted. Our asset generator is trained to disentangle labeled and unlabeled information and therefore designed to excel in this aspect. However, there is no direct metric proposed in the past that can conveniently evaluate this capability quantitatively: existing face embedding networks for facial identification take all semantic information into consideration, since age, gender and ethnicity are posed as part of one’s identity in these models.
We therefore devise a similarity metric based on linear fitting to evaluate identity preservation of our model in gender and age editing. With the existing face embedding network OpenFace (Amos et al., 2016), we synthesize a large number of random pairs of images, where the two codes for each pair are identical except the controlled attributes. Then, the difference vectors of the embeddings are computed from the two images. The large set of difference vectors are used for linear fitting, where the fitted line should represent the direction corresponding to the controlled attributes (e.g. gender, age) in the embedding space. Note that we exclude ethnicity from this evaluation since its dimensionality is unknown.
The gender-and-age-independent identity similarity can therefore be obtained by projecting the face embeddings onto the complement of the subspace spanned by the estimated gender and age directions. The similarity of two images can be measured by the cosine similarity. We therefore compare the identity similarity of pairs of images where gender and age attributes are edited by our models against random pairs of face images. The results are shown in Table 5, where we find that our model supports attribute editing that maintains a high similarity in identity under our metric. If we change just the age and/or gender, the result should have high similarity to the original, while two random images should have lower similarity.
| change gender | change age | random pair |
Evaluation on the Two-Step Disentanglement Training.
To evaluate the effectiveness of the disentanglement training, we compare our two-stage training design to a typical conditional GAN model as the baseline approach. Examples are shown in Figure 18. We find that our generator can better reproduce unlabeled features (e.g. facial identity): In column 1, the result produced by the baseline GAN has flatter eyebrows; in columns 2 and 4 the baseline result displays less facial hair; in column 3 the baseline result appears more masculine; in column 5 the baseline result shows less wrinkles.
4.2.4. Application